### v1

# D-tensor paraproducts and its caricatures

Abstract
We generalize the $2$-tensor paraproduct decomposition result of [1] to $d$-tensors and expound on its caricatures. In particular, we show that for $A\in C^{d},f\in\Lambda^{\alpha}([0,1]^{d})$, $A(f)$ has an approximation $\tilde{A}(f)=(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f))$ for a fixed sequence of scales, $\mathbf{j}=(j_{1},j_{2},\ldots,j_{d})$, and the series $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f)$ is a taylor expansion of $A$. Additionally, we show the sequence of operators $(\tilde{\mathbf{v}}^{1}(f),\tilde{\mathbf{v}}^{2}(f),\ldots,\tilde{\mathbf{v}}^{d}(f))$ form a basis for the subspace comprised of linear functionals of the form $A(f)$, and the residual, $\Delta(A,f)=\tilde{A}(f)-A(f)\in\Lambda^{2\alpha}([0,1]^{d})$. Consequentially, we show one can obtain a Calderon-Zygmund type decomposition, $\Delta(A,f)_{\lambda}+\tilde{A}(f)_{\lambda}$ with the obtained paraproduct decomposition $A(f)=\Delta(A,f)+\tilde{A}(f)$. Our theoretical findings are supported by computational examples for d=2,3.
1 Introduction
We generalize the 2-tensor paraproduct decomposition established in [1] to $d$-tensors. A version of J.M. Bony’s celebrated result [2] was formulated in [1] for smooth compositions of mixed-holder functions which appear in a variety of matrix processing tasks [3, 4]. We extend this result for Holder functions of the form $A(f),f:[0,1]^{d}\to\mathbb{R},A\in C^{d}$ to obtain the paraproduct decomposition, $A(f)=\tilde{A}(f)+\Delta(A,f)$. Examination of the organs of $\tilde{A}(f)+\Delta(A,f)$ leads to several caricatures which are highlighted herein. Namely, in higher dimensions we show such a decomposition is a taylor expansion of $A$ on the image of $f$, the operators in the taylor expansion of $f$ form a basis for the subspace comprised of functions of the form $A(f)$, and the paraproduct decompostion, $A(f)=\tilde{A}(f)+\Delta(A,f)$, can readily be modified to obtain the Calderon-Zygmund decomposition, $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$.
Our theoretical findings are supported by computational examples of commonly encountered mathematical objects for $d=2,3$. For $d=2$ we consider the potential kernel, $A_{1}(f)=\frac{1}{f(x,y)}$ and the heat kernel $A_{2}(f)=e^{-\frac{f(x,y)}{\epsilon}}$ where $f(x,y)=\lVert x-y\rVert_{l_{2}}$. (we first organize $f(x,y)$ to be $\Lambda^{\alpha}([0,1]^{2})$ using the algorithm from [5]). Next we consider a cone singularity in the complex plane modulated by a graphic equalizer: $A_{3}(f)=\text{sin}(\frac{j_{x}+j_{iy}}{2}f(z))$. Subsequent processing of the low and high frequency components separately with Haar for $A_{1}(f),A_{2}(f)$, for example, leads to tensor compression suggesting similar algorithms to [6] could be explored for $d=2,3$ with such a decomposition. Furthermore, we show the time-varying singularity in the complex plane ($A_{3}(f)$) can be detected in the presence of the graphic equalizer. First, we restate the decomposition theorem for $2$-tensors:
Theorem 1.
Suppose $A\in C^{2}$, $f\in\Lambda^{\alpha}([0,1]^{2}),0<\alpha<\frac{1}{2}$, then we obtain the following discrete multiscale tensor decomposition of $A(f)$:
$A(f)=\sum_{j^{\prime}=0^{\prime}}^{N^{\prime}}\sum_{j=0}^{N}A^{\prime}(P^{j}P^{\prime j^{\prime}}(f))[P^{j+1}P^{\prime j^{\prime}+1}(f)-P^{j}P^{\prime j^{\prime}+1}(f)-P^{j+1}P^{\prime j^{\prime}}(f)+P^{j}P^{\prime j^{\prime}}(f)]$
$+A^{\prime\prime}(P^{j}P^{\prime j^{\prime}}(f))[P^{j+1}P^{\prime j^{\prime}}(f)-P^{j}P^{\prime j^{\prime}}(f)][P^{j}P^{\prime j^{\prime}+1}(f)-P^{j}P^{\prime j^{\prime}}(f)]+\Delta_{N,N^{\prime}}(A,f)$ (1)
$A(f)=\tilde{A}(f)+\Delta_{N,N^{\prime}}(A,f)$ (2)
$\Delta_{N,N^{\prime}}(A,f)=A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{2})$ (3)
where $\Delta_{N,N^{\prime}}(A,f)$ is the residual between $A(f)$ and $\tilde{A}(f)$ up to scales $N,N^{\prime}$. $P^{j}P^{\prime j^{\prime}}$ is a convolution operators acting on $f$ with the tensor product of the scaling function, $\phi^{j,j^{\prime}}_{k,k^{\prime}}(x,x^{\prime})$ (here defined to be the characteristic function) at scales $j,j^{\prime}$ as its kernel. $R^{j,j^{\prime}}_{k,k^{\prime}}\coloneq I^{j}_{k}\times I^{j^{\prime}}_{k^{\prime}}$ is a dyadic rectangle constructed from taking the product of dyadic intervals $I^{j}_{k}\coloneq(2^{-j}k,2^{-j}(k+1)],k,j\in\mathbb{N}$ and $I^{j^{\prime}}_{k^{\prime}}\coloneq(2^{-j^{\prime}}k,2^{-j^{\prime}}(k^{\prime}+1)],k^{\prime},j^{\prime}\in\mathbb{N}$ such that $|R^{j,j^{\prime}}_{k,k^{\prime}}|=2^{-(j+j^{\prime})}$
This decomposition utilizes the connection between the finite difference of the scales of the kernels (here the kernels are the scaling functions) of the averaging operators acting on mixed holder functions to the fundamental theorem of calculus. We can use the same connection to establish the result in $d$-dimensions. We begin by defining mixed-Holder regularity for a $d$-dimensional function as well as the $d$-dimensional scaling function:
Definition 1.
The 1D scaling function is defined as:
$\phi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\phi(2^{-j}x-k),j,k\in\mathbb{N}$ (4)
where the scaling function, $\phi(x)$, is the characteristic function:
$\phi(x)\coloneq\begin{cases}1&\text{if }x\in(0,1]\\ 0&\text{if }x\notin(0,1]\end{cases}$ (5)
Definition 2.
The 1D wavelet function is defined as:
$\psi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\psi(2^{-j}x-k),j,k\in\mathbb{N}$ (6)
where the wavelet function, $\phi(x)$ is the Haar function:
$\psi(x)\coloneq\begin{cases}1&\text{if }0<x\leq\frac{1}{2}\\ -1&\text{if }\frac{1}{2}\leq x\leq 1\end{cases}$ (7)
Definition 3.
We define the $d$-dimensional scaling and wavelet functions as the cartesian product of $d$ one-dimensional scaling functions and wavelet functions, respectively.
$\phi(x_{1},x_{2},\ldots,x_{d})\coloneq\phi(x_{1})\times\phi(x_{2})\times\ldots\times\phi(x_{d})$ (8)
$\psi(x_{1},x_{2},\ldots,x_{d})\coloneq\psi(x_{1})\times\psi(x_{2})\times\ldots\times\psi(x_{d})$ (9)
Definition 4.
Let $\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})=f(x^{\prime}_{1},x_{2},\ldots,x_{d})-f(x_{1},x_{2},\ldots,x_{d})$ such that $\delta_{x_{1}}$ is the difference operator in direction $x_{1}$ We say a function is mixed $\alpha$-Holder if the following condition is satisfied:
$\delta_{x_{d}}\delta_{x_{d-1}}\ldots\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})\leq C(|x^{\prime}_{1}-x_{1}|^{\alpha}|x^{\prime}_{2}-x_{2}|^{\alpha}\ldots|x^{\prime}_{d}-x_{d}|^{\alpha}),0<\alpha<\frac{1}{2}$ (10)
We denote the space of functions satisfying this condition as $\Lambda^{\alpha}([0,1]^{d})$
We now define the averaging operator acting on a function supported on a $d$-dimensional box. Let $I_{1}\times I_{2}\times\ldots\times I_{d}=[0,1]^{d}$. Along each interval we construct $j_{i}=0_{i},1_{i},2_{i},\ldots,N_{i}$ scales such that at scale $j$ one has $k_{i}=1_{i},2_{i},\ldots,2^{j_{i}}$ intervals. We denote the $k$th dyadic interval along scale $j$ in direction $i$ for the interval in the ith direction, $I_{i}$, as $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1))$ and $\chi^{j_{i}}_{k_{i}}(x_{i})$ as the characteristic function associated with interval $I^{j_{i}}_{k_{i}}$.
Definition 5.
We define the averaging operator acting on a function supported on a box as:
$P^{j_{1},j_{2},\ldots,j_{d}}(f)\coloneq\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}P^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f)$ (11)
$=\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}\frac{1}{|I^{j_{1}}_{k_{1}}|\times|I^{j_{2}}_{k_{2}}|\times\ldots\times|I^{j_{d}}_{k_{d}}|}(\int_{|I^{j_{1}}_{k_{1}}|}\int_{|I^{j_{2}}_{k_{2}}|}\ldots\int_{|I^{j_{d}}_{k_{d}}|}$ (12)
$\phi(y_{1},y_{2},\ldots,y_{d})f(y_{1},y_{2},\ldots,y_{d})dy_{1}dy_{2},\ldots dy_{d})\chi^{j_{1}}_{k_{1}}(x_{1})\times\chi^{j_{2}}_{k_{2}}(x_{2})\times\ldots\times\chi^{j_{d}}_{k_{d}}(x_{d})$ (13)
Definition 6.
Let the dyadic distance along direction $i$ be defined as:
$d_{i}(x^{\prime}_{i},x_{i})\coloneq\inf_{I_{i}}\{|I_{i}|:x^{\prime}_{i},x_{i}\in I_{i}\}$ (14)
2 d-tensor paraproducts
Theorem 2.
Suppose $A\in C^{d},f\in\Lambda^{\alpha}([0,1]^{d})$, then we obtain the following discrete $d$-tensor paraproduct decomposition for $A(f)$:
${A}(f)=\sum_{j_{1}=0_{1}}^{N_{1}}\sum_{j_{2}=0_{2}}^{N_{2}}\ldots\sum_{j_{d}=0_{d}}^{N_{d}}(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f))+\Delta(A,f)$ (15)
$A(f)=\tilde{A}(f)+\Delta(A,f)$ (16)
where the residual $\Delta(A,f)$ is:
$\Delta(A,f)\coloneq A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{d})$ (17)
$P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is the averaging operator acting on $f$ at scales $(j_{1},j_{2},\ldots,j_{d})$ and $\tilde{\mathbf{v}}^{\alpha}(f)$ is an operator acting on $f$ comprised of wavelets in $\alpha$ directions.
Proof.
We first prove a short lemma which establishes $A(f)$ can be expanded into a telescoping series with respect the mixed difference of the scales of the averaging operator.
Lemma 1.
Let $A\in C^{\infty}$ while $f\in\Lambda^{\alpha}([0,1]^{d})$. Define $Q^{j_{i}}$ as the finite difference operator between scales of the averaging operator:
$Q^{j_{i}}(f)\coloneq\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}}P_{k_{i}}^{{j_{i}}+1}(f)-\sum_{k_{i}=1_{i}}^{2^{j_{i}}+1}P_{k_{i}}^{j_{i}}(f)$
$=\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}+1}_{k_{i}}|}(\int_{|I^{{j_{i}}+1}_{k_{i}}|}\phi^{{j_{i}}+1}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i})\times\chi^{j_{i}+1}_{k_{i}}(x_{i})$
$-\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}(\int_{|I^{{j_{i}}}_{k_{i}}|}\phi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i})\times\chi^{j_{i}}_{k_{i}}(x_{i})$
$=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$ (18)
where the kernel of $Q^{j_{i}}$ is the wavelet function in direction $i$. Then the following identity holds:
$A(f)=\sum_{j_{d}=0_{d}}^{N_{d}}\sum_{j_{2}=0_{2}}^{N_{d-1}}\ldots\sum_{j_{1}=0_{1}}^{N_{1}}A(Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}})(f)$ (19)
if we assume the average of $f$ over the entire interval in any direction is 0.
Proof.
Let $Q^{j_{m}}(f)$ be the finite difference operator between scales of the averaging operator in direction $m$, where $P^{j_{m}}(f)$ denotes an average over $f$ in direction $m$.
$Q^{j_{i}}(f)=P^{j_{i}+1}(f)-P^{j_{i}}(f)$
$=\sum_{k_{i}=0_{i}}^{2^{N_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$ (20)
Let $A({Q^{j_{i}}}(f))=A(P^{j_{i}+1}(f))-A(P^{j_{i}}(f))$ denote the pointwise application of $A$ on the finite difference operator in direction $i$. The result is established by inducting on the directions the finite difference operator is applied to $f$.
$\sum_{j_{m}=0}^{N_{i}}A({Q^{j_{i}}})(f)=\sum_{j_{i}=0}^{N_{i}}A(P^{j_{i}+1}(f))-A(P^{j_{i}}(f))$
$=A(P^{N_{i}+2}(f))-A(P^{0_{i}}(f))$
$=A(f)$ (21)
where the relation in the second line holds by expansion of the telescoping series, and the relation in the third line hold since $A(P^{j_{i}}(f))\to A(f)$ as $j_{i}\to\infty$ and $A(P^{0_{i}}(f))=0$ by assumption. Now assume this holds for $k$ applications of finite difference operators:
$\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}A({Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}})(f)=A(f)$ (22)
We show the relation holds for $k+1$ applications:
$\sum_{j_{k+1}=0}^{N_{k+1}}(\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}A(Q^{j_{k+1}}Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}})(f)$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(Q^{j_{k+1}}((\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}})(f))$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(Q^{j_{k+1}})(f)$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(P^{j_{k+1}+1}(f))-A(P^{j_{k+1}}(f))$
$=A(P^{N_{k+1}+2}(f))-A(P^{0_{k+1}}(f))$
$=A(f)$ (23)
Since we’ve shown that if the relation holds for the pointwise application of $A$ to $k$ sequential applications of the mixed difference operator to $f$ it holds for the pointwise application of $k+1$ sequential applications, the result holds for the pointwise application of an arbitrary number, $d$, sequential applications of the mixed difference operator to $f$. ∎
The fundamental theorem of calculus can be used to connect the identity established in lemma 3 to a quasilinearization of $A(f)$, but to do so we need to define a $d$-dimensional interpolation operator. This leads us to the next lemma.
Lemma 2.
Define $\delta^{t_{i}}(f)=P^{j_{i}}(f)+t_{i}(P^{j_{i}+1}(f)-P^{j_{i}}(f))$ to be an interpolation operator acting at scale $j_{i}$. Then the following relation holds:
$\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{d}=0}^{t_{d}=1}\left.\right|_{t_{d-1}=0}^{t_{d-1}=1}\ldots\left.\right|_{t_{1}=0}^{t_{1}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$ (24)
Proof.
This can be proved by inducting on the sequential applications of the interpolation operators. First we show the base case holds:
$\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\coloneq P^{j_{1}}(f)+t_{1}(P^{j_{1}+1}(f)-P^{j_{1}}(f))\left.\right|_{t_{1}=0}^{t_{1}=1}$
$=(P^{j_{1}+1}(f)-P^{j_{1}}(f)$
$=Q^{j_{1}}(f)$ (25)
Assume the result holds for $k$ sequential applications of interpolation operators:
$\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1}=Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)$ (26)
We can then show the relation holds for $k+1$ sequential applications of interpolation operators:
$\delta^{t_{k+1}}(\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{0}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1})\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}=\delta^{t_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}$
$=P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))+t_{k+1}(P^{j_{k+1}+1}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))-P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)))\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}$
$=(P^{j_{k+1}+1}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))-P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)))$
$=(Q^{j_{k+1}}Q^{j_{k}}\ldots Q^{j_{1}}(f))$ (27)
Since the relation holds for $k+1$ sequential applications of the interpolation operator given the assumption it holds for $k$ sequential applications of the interpolation operator, by induction it holds for $d$ sequential applications of the interpolation operator. Thus, we have:
$\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{d}=0}^{t_{d}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$ (28)
which completes the proof.
Definition 7.
Let $\delta^{t}(f)$ be an interpolation function which interpolates between the scales $(j_{1},j_{2},\ldots,j_{d})$ of the averaging operators with scaling parameters $(t_{1},t_{2},\ldots,t_{d})$ with $0\leq t_{k}\leq 1$ for $k=1,2\ldots,m$:
$\delta^{t}(f)\coloneq\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)$ (29)
such that lemma 2 holds
From the previous two lemmas and the fundamental theorem of calculus $A(f)$ has an identity which is a summation (with respect to the scaling parameters in all directions) over an integration of the power series comprised of the operators $(\mathbf{v}^{1}(f),\mathbf{v}^{2}(f),\ldots,\mathbf{v}^{d}(f))$ with respect to the interpolation parameters in all directions. First, we define the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ (setting each $\mathbf{v}(f)\coloneq\mathbf{v}^{m}$ to simplify notation) which come from differentiating $A(\delta^{t}(f))$. The elements in the sequence $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ are defined such that each power of $\mathbf{v}$ corresponds to the number of independent partial derivatives of $\delta^{t}(f)$ taken in each term of the series use to construct some $\mathbf{v}^{m}\in(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$:
$\mathbf{v}^{1}\coloneq\frac{\partial^{d}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (30)
$\mathbf{v}^{2}\coloneq\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\frac{\partial}{\partial t_{2}}(\delta^{t}(f))\frac{\partial}{\partial t_{1}}(\delta^{t}(f)))+$
$\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{d-1}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (31)
$\vdots$
$\mathbf{v}^{m}\coloneq\frac{\partial^{m}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
$+\frac{\partial^{m-1}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{d-m+1}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-m}\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\ldots+$
$\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{m}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{d-m-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-m-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (32)
$\vdots$
$\mathbf{v}^{d-1}\coloneq\frac{\partial^{2}}{\partial t_{d}\partial t_{d-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-1}\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
$+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (33)
$\mathbf{v}^{d}\coloneq\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (34)
Note that the first, $\mathbf{v}^{1}$, and last, $\mathbf{v}^{d}$, operators have one term while the operators are comprised of a series - each term in the series consisting of the number of partial derivatives corresponding to the power of the operator (i.e. the term $\mathbf{v}^{m}$ has $m$ partial derivatives for each term in its series). The identity for $A(f)$ can compactly be written as:
$A(f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}A(\delta^{t}(f))\mathbf{v}^{1}+A(\delta^{t}(f))\mathbf{v}^{2}+\ldots+$
$A^{m}(\delta^{t}(f))\mathbf{v}^{m}+\ldots+A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}+A^{d}(\delta^{t}(f))\mathbf{v}^{d}dt_{1}dt_{2}\ldots dt_{d}$ (35)
We can now define the approximation, $\tilde{A}(f)$, by replacing the sequence $(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$ with $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ - each element in the latter sequence is comprised with a wavelet coefficient in the direction of the partial derivative for an element in the former sequence.
$\mathbf{\tilde{v}}^{1}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (36)
$\tilde{\mathbf{v}}^{2}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{2}}(f)Q^{j_{1}}(f)+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)+\ldots$
$+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (37)
$\vdots$
$\tilde{\mathbf{v}}^{m}\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)$
$+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{d-m+1}}(f)Q^{j_{d-m}}Q^{j_{m}}(f)\ldots Q^{j_{m-1}}(f)+\ldots+$
$Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{d-m-1}}(f)Q^{j_{d-m-2}}(f)\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (38)
$\vdots$
$\tilde{\mathbf{v}}^{d-1}\coloneq Q^{j_{d}}Q^{j_{d-1}}(f)Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)$
$+\ldots+Q^{j_{d}}(f)Q^{j_{d-1}}(f))\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (39)
$\tilde{\mathbf{v}}^{d}\coloneq Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)\in[0,1]^{d}\to\mathbb{R}$ (40)
Again, note the first, $\tilde{\mathbf{v}}^{1}$, and last, $\tilde{\mathbf{v}}^{d}$, operators have one term while the operators, $(\mathbf{v}^{2},\ldots,\mathbf{v}^{d-1})$, are comprised of a series each with d1d-1 terms. In this case, each term in the series contains the number of wavelet coefficients corresponding to the power of the operator (i.e. the term $\tilde{\mathbf{v}}^{m}$ has $m$ wavelet coefficients for each term in its series). Now the following linear combination of $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ is independent of $t$ and is used to construct the approximation, $\tilde{A}(f)$:
$\tilde{A}(f)\coloneq\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}A(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}+A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2}+\ldots+A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m}$
$+\ldots+A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$
$=\sum_{j_{d}=0_{d}}^{N_{d}}\sum_{j_{d-1}=0_{d-1}}^{N_{d-1}}\ldots\sum_{j_{1}=0_{1}}^{N_{1}}(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha})$ (41)
Computing the residual, $\Delta(A,f)\coloneq A(f)-\tilde{A}(f)$, yields:
[…]
$(A^{\prime\prime}(\delta^{t}(f))\mathbf{v}^{2}-A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2})+\ldots+(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})+\ldots+$
$(A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}-A^{d-1}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d-1})+(A^{d}(\delta^{t}(f))\mathbf{v}^{d}-A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})dt_{1}dt_{2}\ldots dt_{d}$
$=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})dt_{1}dt_{2}\ldots dt_{d}$ (42)
We characterize the regularity of the residual, $\Delta(A,f)$, by estimates on the terms in the integrand. Consider an arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the series, $\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand. Then
$\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
from Holders inequality. Handling the estimates of $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ separately, we have:
$\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$
$=\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$
$\leq\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{1}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\int_{\alpha^{\mathbf{j}}_{\mathbf{k}}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}|A^{m}(y)|dy$
$\leq\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (44)
where $\alpha^{\mathbf{j}}_{\mathbf{k}}$ is the wavelet coefficient at the specified scale and location parameters, the second inequality comes from triangle inequality, the third inequality from the $L^{1}$ estimate on $A^{m}$- which is finite since $A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\in L^{\infty}$ by the maximum principle - and that wavelet coefficients of Holder functions can be characterized by exponential decay in the scale parameters [7, 8]. For $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$,
$\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
$=(d-1)\lVert Q^{j_{1},j_{2},\ldots,j_{d}}\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (45)
where again, the last inequality comes from exponential decay in the wavelet scaling parameter and Holder functions. Combining the estimates on $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ yields:
$\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$ (46)
Now we bound the approximation terms, $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}$, for the mth term in the series under the integral for $\Delta(A,f)$ similarly. By Holder’s inequality,
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ (47)
where
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (48)
where the last inequality holds from the same argument in (44), and
$\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}=\lVert Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}+$
$\lVert Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{d-m+1}}(f)Q^{j_{m-1}}(f)Q^{j_{m}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$+\ldots+\lVert Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{d-m-1}}(f)Q^{j_{d-m-2}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{d}+j_{d-1}+\ldots+j_{m})\alpha}2^{-j_{m-1}\alpha}\ldots 2^{-j_{1}\alpha}$
$+2^{-(j_{d}+j_{d-1}+\ldots+j_{d-m+1})\alpha}2^{-j_{m-1}\alpha}\ldots 2^{-j_{1}\alpha}+\ldots+2^{-j_{d}\alpha}2^{-(j_{d-1}+\ldots+j_{d-m-1})\alpha}2^{-j_{d-m-2}\alpha}\ldots\ldots 2^{-j_{1}\alpha}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (49)
giving us,
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (50)
Combining the inequalities in (46) and (50) gives:
$\left|\lVert(A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (51)
The argument above used to provide estimates for the arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand holds for the terms in the series, $\sum_{m=2}^{d-1}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$. The first and last terms are handled slightly different since $\mathbf{v}^{1}$ and $\mathbf{v}^{d}$ contain 1 term in their series instead of d1d-1 terms. Since $\mathbf{v}^{1}=\tilde{\mathbf{v}}^{1}$, the first term becomes:
$(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}$ (52)
Then the following $L^{1}$ estimate is obtained:
$\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}\rVert_{L^{1}([0,1]^{d})}\leq\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\rVert_{L^{1}([0,1]^{d})}\rVert\tilde{\mathbf{v}}^{1}\lVert_{L^{\infty}([0,1]^{d})}$
$=\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\lVert\delta^{t}(f))-P^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{1}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\int_{\alpha^{\mathbf{j}}_{\mathbf{k}}(I^{j}_{k}\times I^{j}_{k}\times\ldots\times I^{j}_{k})}Q(y)dy\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (53)
by application of Holder’s inequality, the property of $A\in C^{d}$, and the estimates on the wavelet coefficients of Holder functions previously discussed. So by the previous result, the first term in the series under the integrand of $\Delta(A,f)$ has the estimate:
$\left|\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (54)
Now we handle the final term in the series, $(A^{d}(\delta^{t}(f))\mathbf{v}^{d}-A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$, where $\tilde{\mathbf{v}}^{d}$ only has 1 term in its series. Since $\mathbf{v}\neq\tilde{\mathbf{v}}^{d}$, we handle $(A^{d}(\delta^{t}(f))\mathbf{v}^{d}$ and $A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$ separately as we did for the $m$th term; the obtained estimates are the same but the technology used is negligibly different. The $\lVert A^{d}(\delta^{t}(f))\rVert_{L^{1}([0,1])}$ estimate is the same as (45), but $\lVert\mathbf{v}^{d}\rVert_{L}^{\infty}([0,1])$ is slightly different:
$\lVert\mathbf{v}^{d}\rVert_{L^{\infty}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}\mathbf{v}^{d}\rVert_{L^{\infty}([0,1]^{d})}$
$=\lVert Q^{j_{1},j_{2},\ldots,j_{d}}\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (55)
Thus, we again have:
$\lVert A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$ (56)
as in (46). Similar reasoning applies to the $A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$ term. The $\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$ estimate is the same as (48), but for $\lVert\tilde{\mathbf{v}}^{d}\rVert_{L^{\infty}([0,1]^{d})}$ we have:
$\lVert\tilde{\mathbf{v}}^{d}\rVert_{L^{\infty}([0,1]^{d})}=\lVert Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq\sum_{k_{d}=1_{d}}^{2^{j_{d}}}Q^{j_{d}}_{k_{d}}(f)\sum_{k_{d-1}=1_{d-1}}^{2^{j_{d-1}}}Q^{j_{d-1}}_{k_{d-1}}(f)\ldots\sum_{k_{1}=1_{1}}^{2^{j_{1}}}Q^{j_{1}}_{k_{1}}(f)$
$\leq C\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-j_{d}\alpha}\sum_{k_{d-1}=1_{d-1}}^{2^{j_{d-1}}}2^{-j_{d-1}\alpha}\ldots\sum_{k_{1}=1_{1}}^{2^{j_{1}}}2^{-j_{1}\alpha}$
$=C2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$ (57)
Combining (56) and (57), the last term of the series in the integrand of $\Delta(A,f)$ has the estimate:
$\left|\lVert(A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$ (58)
So by (51), (54), and (58) we can express the residual, $\Delta(A,f)$, defined in (42), as a sum of piecewise constant functions:
$\Delta(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
$|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$ (59)
Now we characterize the regularity of $\Delta(A,f)$ with the following lemma:
Lemma 3.
The residual, $\Delta(A,f)=A(f)-\tilde{A}(f)$, defined as:
$\Delta(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
$|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}([0,1]^{d})$ (60)
which is the sum of piecewise constant functions with bounded coefficients, is an element of $\Lambda^{2\alpha}$
Proof.
To show $\Delta(A,f)\in\Lambda^{2\alpha}([0,1]^{d})$, we need to show
$|\Delta(A,f)(\mathbf{x})-\Delta(A,f)(\mathbf{\tilde{x}})|\leq Cd_{1}(x_{1},\tilde{x}_{1})^{2\alpha}d_{2}(x_{2},\tilde{x}_{2})^{2\alpha}\ldots d_{d}(x_{d},\tilde{x}_{d})^{2\alpha}$ (61)
[…]
Similarly, let $\mathcal{L}$ be the set containing combinations of scaling and location parameters such that $\mathbf{x},\tilde{\mathbf{x}}$ are never contained in the same box.
[…]
Then summing over the scaling and location parameters in $\mathcal{U}$ for $\Delta(A,f)$, one has:
$\sum_{(\mathbf{j},\mathbf{l})\in\mathcal{U}}\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{x})-\gamma^{\mathbf{{j}}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{\tilde{x}})=0$
$\leq Cd_{1}(x_{1},\tilde{x}_{1})^{2\alpha}d_{2}(x_{2},\tilde{x}_{2})^{2\alpha}\ldots d_{d}(x_{d},\tilde{x}_{d})^{2\alpha}$ (64)
since $\Delta(A,f)$ is piecewise constant on dyadic boxes (see 59). Now consider the case where $x,x^{\prime}$ are not contained in same dyadic box. Again, summing over the scaling and location parameters in $\mathcal{L}$ for $\Delta(A,f)$ one has:
$\sum_{(\mathbf{j},\mathbf{l})\in\mathcal{L}}\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{x})-\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{\tilde{x}})\leq C\sum_{\mathbf{j}^{(i)}=\mathbf{j}^{(0)}}^{\mathbf{j}^{(N)}}2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(i)})2\alpha}$
$=\sum_{\mathbf{j}^{(i)}=\mathbf{j}^{(0)}}^{\mathbf{j}^{(N)}}2^{-(j^{(i)}_{1}+j^{(i)}_{2}+\ldots+j^{(i)}_{d})2\alpha}$
$\leq C2^{-(j^{(N)}_{1}+j^{(N)}_{2}+\ldots+j^{(N)}_{d})2\alpha}$
$\leq Cd_{1}(x_{1},x^{\prime}_{1})^{2\alpha}d_{1}(x_{2},x^{\prime}_{2})^{2\alpha}\ldots d_{d}(x_{d},x^{\prime}_{d})^{2\alpha}$ (65)
where we order the scales in $\mathcal{L}$ such that $\mathbf{j}^{(0)}\leq\mathbf{j}^{(1)}\leq\ldots\leq\mathbf{j}^{(N)}$ indicates $2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(0)})}\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(1)})}\leq\ldots\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(N)})}$. The first inequality holds from the estimates on $\Delta(A,f)$ obtained in (59), the second inequality from collapsing the geometric series, the third from the assumption on the scaling parameters contained in $\mathcal{L}$. Thus, combining the previous two estimates in (64) and (65) one has:
[…]
$\leq Cd_{1}(x_{1},x^{\prime}_{1})^{2\alpha}d_{1}(x_{2},x^{\prime}_{2})^{2\alpha}\ldots d_{d}(x_{d},x^{\prime}_{d})^{2\alpha}$ (66)
Since the choice of $\mathbf{x},\mathbf{\tilde{x}}$ were arbitrarily chosen, $\Delta(A,f)\in\Lambda^{2\alpha}([0,1]^{d})$
Thus, from lemma 3, we can compute the approximation $\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ to $A(f)$ such that $\Delta(A,f)=A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{d})$.
3 Paraproduct Caricatures
3.1 3-tensor paraproducts and taylor expansion of $\tilde{A}(f)$
Corollary 1.
Suppose $A\in C^{3},f\in\Lambda^{2\alpha}([0,1]^{3})$, then we obtain the following discrete $3$-tensor paraproduct decomposition
$\tilde{A}(f)=\sum_{j_{1}=0_{1}}^{N_{1}}\sum_{j_{2}=0_{2}}^{N_{2}}\sum_{j_{3}=0_{3}}^{N_{3}}A^{\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{1}}(f)Q^{j_{2}}(f)Q^{j_{3}}(f)+$
$+A^{\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))[Q^{j_{2}}[Q^{j_{1}}(f)Q^{j_{3}}(f)]+Q^{j_{3}}[Q^{j_{2}}(f)Q^{j_{1}}(f)]]$
$+A^{\prime\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{3}}Q^{j_{2}}Q^{j_{1}}(f)$ (67)
Proof.
The proof follows from the same techniques used to prove the general $d$-tensor case by setting $d=3$
Corollary 2.
The approximation, $\tilde{A}(f)$, is real analytic on the image of $f$ and consequently can be expanded into a taylor series expansion for fixed scales $(j_{1},j_{2},\ldots,j_{d})$
$\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ (68)
where $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is the derivative of A of order $\alpha$ evaluated at the average of $f$ and $\tilde{\mathbf{v}}^{\alpha}$ is an operator acting on $f$ containing wavelet coefficients in $\alpha$ directions.
Proof.
As any real analytic function can be represented by its taylor series, it suffices to show $A$ is real analytic on $f$. Since $\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ by Theorem 2 we argue $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ has a power series representation. $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is a constant coefficient for $\alpha=1,2,\ldots,d$ and since $\tilde{\mathbf{v}}^{\alpha}$ represents $\alpha$ products of wavelet coefficients, $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ has a power series representation; therefore $A$ is real analytic on $f$ and has the taylor expansion $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ where $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ are the coefficients in the expansion which are derivatives of $A$. ∎
Corollary 3.
The sequence of operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ form a basis for the subspace of linear functionals of the form $U=\{g:g=A(f),f\in\Lambda^{\alpha}([0,1]^{d}),A\in C^{\infty}\}\subset V=\{h:h:[0,1]^{d}\to\mathbb{R}\}$ for a fixed sequences of scales $(j_{1},j_{2},\ldots,j_{d})$
Proof.
$V$ is a vectors space since for $h_{i},h_{j}\in V$ $h_{i}+h_{g}\in V$ and $\alpha h_{i}\in V$ for $\alpha\in\mathbb{R}$. $U\subset V$ is a subspace of $V$ by the same reasoning. Since $A(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ by Theorem 2, the sequence of operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ span $U$. Suppose $A(f)=0$ such that $0=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$. Then the operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ are linearly independent since $A^{1}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=A^{2}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=\cdots=A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=0$ since the derivative of the 0 function is 0. Thus, since $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ span $U$ and are linearly independent, they form a basis for the subspace of linear functionals of the form: $U=\{g:g=A(f),f\in\Lambda^{\alpha}([0,1]^{d}),A\in C^{\infty}\}\subset V=\{h:h:[0,1]^{d}\to\mathbb{R}\}$
3.2 Calderon-Zygmund Decomposition
Theorem 3.
Consider the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$ as defined in Theorem 2. If we assume the average of $f$ with respect to each direction is 0, the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$, can be modified to obtain the following Calderon-Zygmund decomposition $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$, where $\tilde{A}(f)_{\lambda}$ is the bad part of $A(f)_{\lambda}$ and $\Delta(A,f)_{\lambda}$ is the good part of $A(f)$.
Proof.
In the interest of intelligibility, the analysis is carried out, without loss of generality, for a fixed sequence of scales $(j_{1},j_{2},\ldots,j_{d})$ and fixed order, $\alpha=m$, for the approximation term. First we provide a more explicit definition of the projection operator on the wavelet basis (from (18)) in terms of a series expansion of the wavelet coefficients supported on dyadic intervals, $k_{i}=1_{i},\ldots,2^{j_{i}}$, in direction i.
$Q^{j_{i}}(f):=\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}+1}P_{\tilde{k}_{i}}^{{j_{i}}+1}(f)-\sum_{k_{i}=1_{i}}^{2^{j_{i}}}P_{k_{i}}^{j_{i}}(f)$
$=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}Q^{j_{i}}_{k_{i}}(f)$ (69)
Similarly, a more explicit definition of $\tilde{\mathbf{v}}^{m}$, originally defined in (38), in terms of the wavelet coefficients defined on $d$-dimensional dyadic boxes is provided:
$\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)\coloneq Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{m}}^{j_{m}}(f)Q_{k_{m-1}}^{j_{m-1}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$
$+Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{d-m+1}}^{j_{d-m+1}}(f)Q_{k_{d-m}}^{j_{d-m}}Q_{k_{m}}^{j_{m}}(f)\ldots Q_{k_{m-1}}^{j_{m-1}}(f)+\ldots+$
$Q_{k_{d}}^{j_{d}}(f)Q_{k_{d-1}}^{j_{d-1}}Q_{k_{d-2}}^{j_{d-2}}\ldots Q_{k_{d-m-1}}^{j_{d-m-1}}(f)Q_{k_{d-m-2}}^{j_{d-m-2}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$ (70)
$\tilde{\mathbf{v}}^{m}\coloneq\sum_{k_{1}=1_{1},k_{2}=1_{2},\ldots,k_{d}=1_{d}}^{2^{j_{1}},2^{j_{2}},\ldots,2^{j_{d}}}\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)$
$=\sum_{k_{1}=1_{1},k_{2}=1_{2},\ldots,k_{d}=1_{d}}^{2^{j_{1}},2^{j_{2}},\ldots,2^{j_{d}}}Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{m}}^{j_{m}}(f)Q_{k_{m-1}}^{j_{m-1}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$
$+Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{d-m+1}}^{j_{d-m+1}}(f)Q_{k_{d-m}}^{j_{d-m}}Q_{k_{m}}^{j_{m}}(f)\ldots Q_{k_{m-1}}^{j_{m-1}}(f)+\ldots+$
$Q_{k_{d}}^{j_{d}}(f)Q_{k_{d-1}}^{j_{d-1}}Q_{k_{d-2}}^{j_{d-2}}\ldots Q_{k_{d-m-1}}^{j_{d-m-1}}(f)Q_{k_{d-m-2}}^{j_{d-m-2}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$ (71)
Its also profitable to define the set of all dyadic boxes at scales $(j_{1},j_{2},\ldots,j_{d})$:
$\mathcal{D}_{\lambda}\coloneq\{B_{k_{1},k_{2},\ldots,k_{d}}:B_{k_{1},k_{2},\ldots,k_{d}}\coloneq I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}\hskip 5.69046pt\forall\hskip 5.69046ptk_{1}=1_{1}\ldots 2^{j_{1}},k_{2}=1_{2}\ldots 2^{j_{2}},\ldots,$
$k_{d}=1_{d}\ldots 2^{j_{d}},B_{jk_{1},k_{2},\ldots,k_{d}}\coloneq\text{supp}(A^{m}(P_{k_{1},k_{2},\ldots,k_{d}}^{j_{1},j_{2},\ldots,j_{d}}(f))\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)),$
$|A^{m}(P_{k_{1},k_{2},\ldots,k_{d}}^{j_{1},j_{2},\ldots,j_{d}}(f))\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)|>\lambda\coloneq 2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}\}$ (72)
Let each $B_{k_{1},k_{2},\ldots,k_{d}}\in D_{\lambda}$ be indexed as $i=1,2,\ldots,|D_{\lambda}|$ such that each $B_{k_{1},k_{2},\ldots,k_{d}}\in D_{\lambda}$ corresponds to some $B_{i}\in D_{\lambda}$. Similarly, let $B_{k_{1},k_{2},\ldots,k_{d}}\in[0,1]^{d}\setminus D_{\lambda}$ be indexed by $p=1,2,\ldots,|B_{p}\in[0,1]^{d}\setminus D_{\lambda}|$ such that each $B_{k_{1},k_{2},\ldots,k_{d}}\in[0,1]^{d}\setminus D_{\lambda}$ corresponds to some $B_{p}\in[0,1]^{d}\setminus D_{\lambda}$. The set $D_{\lambda}$ contains all dyadic boxers where the $m$th term in the approximation larger than $\lambda$ for a fixed sequence of scales, $(j_{1},j_{2},\ldots,j_{d})$. We now define the purported Calderon-Zygmund decomposition in terms of the paraproduct decomposition by restricting the supports of the good and bad functions :
$A(f)=\tilde{A}(f)+\Delta(A,f)$
$=\sum_{B_{i}\in D_{\lambda}}\tilde{A}(f)+\Delta(A,f)+\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\tilde{A}(f)+\Delta(A,f)$
$=(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))+(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))$
$=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$ (73)
where we have defined the bad function as $\tilde{A}(f)_{\lambda}\coloneq(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))$ and the good function as $\Delta(A,f)_{\lambda}\coloneq(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))$
From [9], if $A(f)\in L^{1}(\mathbb{R}^{d})$ and the following conditions are satisfied,
  1. 1.
    $A(f)=(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))+\Delta(A,f)_{\lambda}$
  2. 2.
    $\lVert\Delta(A,f)_{\lambda}\rVert_{L^{\infty}([0,1]^{d})}\leq C\alpha,\lVert\Delta(A,f)_{\lambda}\rVert_{L^{1}([0,1]^{d})}\leq C\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$
  3. 3.
    $\text{supp}(\tilde{A}_{i}(f)))\subset B_{i}$ , $\int\tilde{A}_{i}(f))=0\forall i=1,2,\ldots|D_{\lambda}|$, $\sum_{B_{i}\in D_{\lambda}}\lVert\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}\leq C\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$
  4. 4.
    $\sum_{B_{i}\in D_{\lambda}}m(B_{i})\leq\frac{C}{\alpha}\lVert A(f)\rVert_{L^{1}}$
where $C>0,\alpha>0$, then the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$ can be represented as a Calderon-Zygmund decomposition:  $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$.
Since $A(f)$ is continuous with compact support, $A(f)\in L^{1}([0,1]^{d})$, satisfying the requisite condition on $A(f)$.Condition (1) follows immediately from Theorem, (2) and (73). The first part of condition (2) holds since:
$\lVert\Delta(A,f)_{\lambda}\rVert_{L^{\infty}([0,1]^{d})}=\lVert\displaystyle\sum_{B_{i}\in D_{\lambda}}\beta_{i}\chi_{B_{i}}(x_{1},x_{2},\ldots,x_{d})+\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d}))2\alpha}$ (74)
by the estimates on the coefficients, $\beta_{i}$, defined in (59) and by construction of $\tilde{A}_{i}(f)$. The second part of condition (2) holds since
$\lVert\Delta(A,f)_{\lambda}\rVert_{L^{1}([0,1]^{d})}=\int_{[0,1]^{d}}|(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))|dx$
$=\int_{[0,1]^{d}}|\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\tilde{A}(f)+\Delta(A,f)|dx$
$=\int_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}|\tilde{A}(f)+\Delta(A,f)|dx$
$\leq C\int_{[0,1]^{d}}|\tilde{A}(f)+\Delta(A,f)|dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (75)
The first part of condition (3) by construction of the set $D_{\lambda}$ in (72), the second part of condition (3) by the assumption that the average of $f$ in each direction is 0, and finally the last part of condition 3 holds since:
$\displaystyle\sum_{B_{i}\in D_{\lambda}}\lVert\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}=\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{[0,1]^{d}}|\tilde{A}_{i}(f)|dx$
$=\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{B_{i}}|\tilde{A}_{i}(f)|dx$
$\leq C(\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{B_{i}}|\tilde{A}_{i}(f)|dx+\displaystyle\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\int_{B_{p}}|\tilde{A}_{i}(f)|)dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (76)
where the last equality holds since $B_{i}\cap B_{p}=\emptyset$, and the inequality holds by the definition of $A(f)$. Finally condition (4) is met since:
$\sum_{B_{i}\in D_{\lambda}}m(B_{i})\leq C\int_{[0,1]^{d}}|A(f)|dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (77)
which holds since $D_{\lambda}\subset[0,1]^{d}$, thus completing the proof.
4 Computational Examples
4.1 Quasilinearization of Heat and Potential Kernels (d=2)
As described in the previous section, the paraproduct and Calderon-Zygmund decompositions are connected; we illustrate this with tensor compression examples of the heat and potential kernels. The approximation captures the ”bad” part, (i.e. the part high frequency content) and has small support, while the residual captures the ”good” part (i.e. the low frequency content).
We process the original potential and heat kernels directly with the standard tensor haar basis, along with the paraproduct decomposition of both kernels. To do this, we randomly generated $\{x_{i}\}_{i=1}^{512}$ source and $\{y_{i}\}_{i=1}^{512}$ target points with each $y_{i},x_{i}\in[-1,1]^{3}$.We take $f(x,y)=\lVert x-y\rVert_{l_{2}}$. We organize $f$ to be $\alpha$-Holder using the questionnaire algorithm. We define $A_{1}(f)=\frac{1}{\lVert x-y\rVert_{l_{2}}}$ as the potential kernel and $A_{2}(f)=e^{-\frac{\lVert x-y\rVert_{l_{2}}}{\epsilon}}$ with $\epsilon=1$ as the heat kernel.
Figure 1: Points generated for heat and potential kernels
Figure 2: top row, left to right: $f,\tilde{A}_{2}(f),\Delta(A_{2},f)$ for heat kernel; bottom row, left to right: $f,\tilde{A_{1}}(f),\Delta(A_{1},f)$ for potential kernel
In Figure 2 we depict the original, approximation, and residual matrices of the heat and potential kernels. We process all of the matrices with the tensor Haar basis and plot the expansion coefficients in Figure 3 below.
Figure 3: (left) Heat kernel (right) Potential kernel expansion coefficients for original, approximation, and residual
In Figure 3 we plot the expansion coefficients for both the heat and potential kernels. In both cases, the matrices of the original functions, $A_{1}(f)$ and $A_{2}(f)$ exhibit slower decay than the approximation and residual matrices from the decomposition. This is quantified in the table below.
Kernels $|$ basis vectors 2000 6000 10000
Original heat kernel, $A_{1}(f)$ 36 % 72 % 91 %
Heat kernel approximation, $\tilde{A_{1}}(f)$ 80 % 98 % 99 %
Heat kernel residual $\Delta(A_{1},f)$ 66 % 91 % 97 %
Original potential kernel, $A_{2}(f)$ 40 % 75 % 92 %
Potential kernel approximation, $\tilde{A_{2}}(f)$ 80 % 98 % 99 %
Potential kernel residual, $\Delta(A_{2},f)$ 55 % 83 % 95 %
Table 1: Percentage of $l_{1}$ entropy captured by top $k\in\{{2000,6000,10000}\}$ basis vectors
Figure 3 and Table 1 demonstrate how the paraproduct decomposition connects to the Calderon-Zygmund decomposition; the high frequency content captured by the approximation has small support and can consequently be captured by only a few tensor basis vectors. The low frequency content captured by the residual has higher regularity ($2\alpha$-Holder vs $\alpha$-Holder) and consequently exhibits faster decay when processed by the tensor wavelet basis. This suggest downstream applications for construction of fast algorithms such as [6].
4.2 Cone singularity in complex plane (d=3)
For the 3D example, we consider a time-varying singularity in the complex plane:
$f(z)=\begin{cases}(1-|z|)^{\alpha}&\text{if }|z|<1\\ (1-\frac{1}{|z|})^{\alpha}&\text{if }\frac{1}{|z|}<1\end{cases}$ (78)
where $|z|=\sqrt{x^{2}+iy^{2}+t^{2}}$ is the usual modulus in the complex plane with a time-dependent parameter and $x,iy\in[-1,1],t\in[0,1]$. $N=256$ equispaced points are sampled in the x and y directions, 128 points are sampled in the time direction. Here, $A(f)$ is the same time-dependent graphic equalization function described in [1], where the amplitude of $f$ is differentially modulated at different frequency bands given by the scales, $j_{x}$ and $j_{iy}$, of the real and imaginary components:
$A_{3}(f)=sin((\frac{j_{x}+j_{iy}}{2}<f(z),\psi^{j_{x},j_{iy}^{\prime}}(x,iy)>\psi^{j_{x},j_{iy}^{\prime}}(x,iy))$ (79)
Figure 4: Cone singularity in complex plane: (a - c) $f,A(f),\tilde{A}(f)$ at t=15 (d- f) $f,A(f),\tilde{A}(f)$ at t=70, (g-i) $f,A(f),\tilde{A}(f)$ at t=115.
The left column of Figure 4 shows the original singularity progressing towards the center, the middle column the nonlinear distortion, the right column the approximation. The rows represent discrete time points, $t=\{15,70,115\}$. Qualitatively, one can see the approximation recovers the ring singularity which is conspicuously distorted by $A(f)$.
We process all the figures in Figure 4 with the tensor haar basis. The red squares are the support of the tensor haar functions with the largest coefficients. This exemplifies the ability of the decomposition to recover the singularity in the presence of smooth non-linear distortions - the singularity is detected as expected for the original function, $f$, goes undetected for the non-linear distortion in the middle column, $A(f)$, and is recovered again by the approximation, $\tilde{A}(f)$ in the right column.
5 Acknowledgments
The author wishes to thank Ronald R. Coifman for useful discussions.