### v2

# D-tensor paraproducts and its caricatures

Abstract.
We generalize the 22-tensor paraproduct decomposition result of [2] to $d$-tensors. In particular, we show that for $A\in C^{d}(\mathbb{R}),f\in\Lambda_{\alpha}([0,1]^{d})$, $A(f)$ can be approximated by $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=(\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f))$ with the residual $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\tilde{A}_{(N_{i})_{i=1}^{d}}(f)-A(f)\in\Lambda_{2\alpha}([0,1]^{d})$. Our theoretical findings are supported by a computational example for d=3.
1. Introduction
We generalize the 2-tensor paraproduct decomposition developed in [2], which built on Bony’s seminal work, [1], to $d$-tensors for the situation $(f\in\Lambda_{\alpha}([0,1]^{d}),A\in C^{d}(\mathbb{R}))$; in particular, we prove the following theorem:
Theorem $1.$
Suppose $A\in C^{d}(\mathbb{R})$, $f\in\Lambda_{\alpha}([0,1]^{d}),0<\alpha<\frac{1}{2}$, then for the operator $T:f\to A(f)$ we can approximate $A(f)$ with
(1) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f)\in\Lambda_{\alpha}([0,1]^{d})$
such that the multiscale tensor paraproduct transforms $T:f\to A(f)$ to
(2) $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}:f\to\tilde{A}_{(N_{i})_{i=0}^{d}}(f)+\Delta_{(N_{i})_{i=0}^{d}}(A,f)$
where $\Delta_{(N_{i})_{i=0}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{2})$ is the residual containing twice the regularity of $f$.
Furthermore, examinations of the organs of $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}$ leads to caricatures highlighted herein - namely, that (a) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ is a taylor expansion of AA in terms of the derivatives of $A$ and scaling parameters of the wavelet operators and (b) the 3-tensor paraproduct decomposition.
2. Acknowledgments
The author wishes to thank Ronald R. Coifman for useful discussions.
3. Preliminaries
Included in this section are the relevant definitions needed to build $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}$:
Definition 3.1.
Dyadic distance along direction $i$:
(3) $d_{d}(x^{\prime}_{i},x_{i})=\inf_{j_{i},k_{i}}\begin{cases}|I^{j_{i}}_{k_{i}}|&x^{\prime}_{i},x_{i}\in I^{j_{i}}_{k_{i}}\\ 0,&x^{\prime}_{i}=x_{i}\end{cases}$
where $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1)]$ is the dyadic interval at scale $j_{i}$ at location $k_{i}$.
Definition 3.2.
The 1D scaling and wavelet functions have the usual definitions:
(4) $\phi(x)\coloneq\begin{cases}1&\text{if }x\in(0,1]\\ 0&\text{if }x\notin(0,1]\end{cases}\qquad\psi(x)\coloneq\begin{cases}1&\text{if }0<x\leq\frac{1}{2}\\ -1&\text{if }\frac{1}{2}<x\leq 1\end{cases}$
(5) $\phi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\phi(2^{-j}x-k),j,k\in\mathbb{N}\qquad\psi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\psi(2^{-j}x-k),j,k\in\mathbb{N}$
Definition $3.$
We define the $d$-dimensional scaling and wavelet functions as the cartesian product of $d$ one-dimensional scaling functions and wavelet functions, respectively.
(6) $\phi^{j}_{k}(x_{1},x_{2},\ldots,x_{d})\coloneq\phi^{j}_{k}(x_{1})\times\phi^{j}_{k}(x_{2})\times\ldots\times\phi^{j}_{k}(x_{d})$
(7) $\psi^{j}_{k}(x_{1},x_{2},\ldots,x_{d})\coloneq\psi^{j}_{k}(x_{1})\times\psi^{j}_{k}(x_{2})\times\ldots\times\psi^{j}_{k}(x_{d})$
Definition 3.4.
Let $\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})=f(x^{\prime}_{1},x_{2},\ldots,x_{d})-f(x_{1},x_{2},\ldots,x_{d})$ such that $\delta_{x_{1}}$ is the difference operator in direction $x_{1}$ We say a function is mixed $\alpha$-Hölder if the following conditions are satisfied:
(8) $\delta_{x_{d}}\delta_{x_{d-1}}\ldots\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})\leq C(d_{d}(x^{\prime}_{1},x_{1})^{\alpha}d_{d}(x^{\prime}_{2},x_{2})^{\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{\alpha}),0<\alpha<\frac{1}{2}$
(9) $\delta_{x_{i}}f(x_{1},x_{2},\ldots,x_{d})\leq Cd_{d}(x^{\prime}_{i},x_{i})^{\alpha}\hskip 5.69046pt,\hskip 5.69046pti=1,\ldots,d$
for $C,\alpha>0$. We denote the space of functions satisfying this condition as $\Lambda_{\alpha}([0,1]^{d})$
We now define the averaging operator acting on a function supported on a $d$-dimensional box. Let $I_{1}\times I_{2}\times\ldots\times I_{d}=[0,1]^{d}$. Along each direction, $i$, we construct $j_{i}=0,1,2,\ldots,N$ scales such that at the $j_{i}$th scale one has $k_{i}=1,2,\ldots,2^{j}$ intervals. Let the dyadic interval along direction $i$ be denoted by: $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1)]$ and let $\chi^{j_{i}}_{k_{i}}(x_{i})$ as the characteristic function associated with interval $I^{j_{i}}_{k_{i}}$.
Definition 3.5.
Let the averaging operator, $P^{j_{1},j_{2},\ldots,j_{d}}$, be understood with the relation:
$P^{j_{1},j_{2},\ldots,j_{d}}(f)\coloneq\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}P^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f)$
$=\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}\frac{1}{|I^{j_{1}}_{k_{1}}|\times|I^{j_{2}}_{k_{2}}|\times\ldots\times|I^{j_{d}}_{k_{d}}|}(\int_{|I^{j_{1}}_{k_{1}}|}\int_{|I^{j_{2}}_{k_{2}}|}\ldots\int_{|I^{j_{d}}_{k_{d}}|}$
(10) $\phi(y_{1},y_{2},\ldots,y_{d})f(y_{1},y_{2},\ldots,y_{d})dy_{1}dy_{2},\ldots dy_{d})\chi^{j_{1}}_{k_{1}}(x_{1})\times\chi^{j_{2}}_{k_{2}}(x_{2})\times\ldots\times\chi^{j_{d}}_{k_{d}}(x_{d})$
4. d-tensor paraproducts
Included in this section is a statement and proof of our main result:
Theorem 4.1.
Suppose $A\in C^{d}(\mathbb{R})$, $f\in\Lambda_{\alpha}([0,1]^{d}),0<\alpha<\frac{1}{2}$, then for the operator $T:f\to A(f)$ we can approximate $A(f)$ with
(11) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f)\in\Lambda_{\alpha}([0,1]^{d})$
such that the multiscale tensor paraproduct transforms $T:f\to A(f)$ to
(12) $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}:f\to\tilde{A}_{(N_{i})_{i=0}^{d}}(f)+\Delta_{(N_{i})_{i=0}^{d}}(A,f)$
where $\Delta_{(N_{i})_{i=0}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{2})$ is the residual containing twice the regularity of $f$.
Proof.
We first prove a short lemma which establishes $A(f)$ can be expanded into a telescoping series with respect the mixed difference of the scales of the averaging operator.
Lemma 4.2.
Let $A\in C^{d}(\mathbb{R})$ and $f\in\Lambda_{\alpha}([0,1]^{d})$. Define $Q^{j_{i}}$ as the finite difference operator between scales of the averaging operator:
$Q^{j_{i}}(f)\coloneq\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}}P_{k_{i}}^{{j_{i}}+1}(f)-\sum_{k_{i}=1_{i}}^{2^{j_{i}}+1}P_{k_{i}}^{j_{i}}(f)$
$=\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}+1}_{k_{i}}|}(\int_{|I^{{j_{i}}+1}_{k_{i}}|}\phi^{{j_{i}}+1}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i})\times\chi^{j_{i}+1}_{k_{i}}(x_{i})$
$-\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}(\int_{|I^{{j_{i}}}_{k_{i}}|}\phi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i})\times\chi^{j_{i}}_{k_{i}}(x_{i})$
(13) $=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$
where the kernel of $Q^{j_{i}}$ is the wavelet function in direction $i$. Then the following identity holds:
(14) $A(f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}A(Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}})(f)$
if we assume the average of $f$ over the entire interval in any direction is 0.
Proof.
Let $Q^{j_{m}}(f)$ be the finite difference operator between scales of the averaging operator in direction $m$, where $P^{j_{m}}(f)$ denotes an average over $f$ in direction $m$.
$Q^{j_{i}}(f)=P^{j_{i}+1}(f)-P^{j_{i}}(f)$
(15) $=\sum_{k_{i}=0_{i}}^{2^{N_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$
Let $A({Q^{j_{i}}}(f))=A(P^{j_{i}+1}(f))-A(P^{j_{i}}(f))$ denote the pointwise application of $A$ on the finite difference operator in direction $i$. The result is established by inducting on the directions the finite difference operator is applied to $f$.
$\sum_{j_{m}=0}^{N_{i}}A({Q^{j_{i}}})(f)=\sum_{j_{i}=0}^{N_{i}}A(P^{j_{i}+1}(f))-A(P^{j_{i}}(f))$
$=A(P^{N_{i}+2}(f))-A(P^{0_{i}}(f))$
(16) $=A(f)$
where the relation in the second line holds by expansion of the telescoping series, and the relation in the third line hold since $A(P^{j_{i}}(f))\to A(f)$ as $j_{i}\to\infty$ and $A(P^{0_{i}}(f))=0$ by assumption. Now assume this holds for $k$ applications of finite difference operators:
(17) $\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}A({Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}})(f)=A(f)$
We show the relation holds for $k+1$ applications:
$\sum_{j_{k+1}=0}^{N_{k+1}}(\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}A(Q^{j_{k+1}}Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}})(f)$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(Q^{j_{k+1}}((\sum_{j_{k}=0}^{N_{k}}\sum_{j_{k-1}=0}^{N_{k-1}}\ldots\sum_{j_{0}=0}^{N_{0}}Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}})(f))$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(Q^{j_{k+1}})(f)$
$=\sum_{j_{k+1}=0}^{N_{k+1}}A(P^{j_{k+1}+1}(f))-A(P^{j_{k+1}}(f))$
$=A(P^{N_{k+1}+2}(f))-A(P^{0_{k+1}}(f))$
(18) $=A(f)$
Since we’ve shown that if the relation holds for the pointwise application of $A$ to $k$ sequential applications of the mixed difference operator to $f$ it holds for the pointwise application of $k+1$ sequential applications, the result holds for the pointwise application of an arbitrary number, $d$, sequential applications of the mixed difference operator to $f$. ∎
The fundamental theorem of calculus can be used to connect the identity established in lemma 4.5 to a quasilinearization of $A(f)$, but to do so we need to define a $d$-dimensional interpolation operator. This leads us to the next lemma.
Lemma 4.3.
Define $\delta^{t_{i}}(f)=P^{j_{i}}(f)+t_{i}(P^{j_{i}+1}(f)-P^{j_{i}}(f))$ to be an interpolation operator acting at scale $j_{i}$. Then the following relation holds:
(19) $\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{d}=0}^{t_{d}=1}\left.\right|_{t_{d-1}=0}^{t_{d-1}=1}\ldots\left.\right|_{t_{1}=0}^{t_{1}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
Proof.
This can be proved by inducting on the sequential applications of the interpolation operators. First we show the base case holds:
$\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\coloneq P^{j_{1}}(f)+t_{1}(P^{j_{1}+1}(f)-P^{j_{1}}(f))\left.\right|_{t_{1}=0}^{t_{1}=1}$
$=(P^{j_{1}+1}(f)-P^{j_{1}}(f)$
(20) $=Q^{j_{1}}(f)$
Assume the result holds for $k$ sequential applications of interpolation operators:
(21) $\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1}=Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)$
We can then show the relation holds for $k+1$ sequential applications of interpolation operators:
$\delta^{t_{k+1}}(\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{0}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1})\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}=\delta^{t_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}$
$=P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))+t_{k+1}(P^{j_{k+1}+1}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))-P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)))\left.\right|_{t_{k+1}=0}^{t_{k+1}=1}$
$=(P^{j_{k+1}+1}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f))-P^{j_{k+1}}(Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)))$
(22) $=(Q^{j_{k+1}}Q^{j_{k}}\ldots Q^{j_{1}}(f))$
Since the relation holds for $k+1$ sequential applications of the interpolation operator given the assumption it holds for $k$ sequential applications of the interpolation operator, by induction it holds for $d$ sequential applications of the interpolation operator. Thus, we have:
(23) $\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{d}=0}^{t_{d}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
which completes the proof.
Definition 4.4.
Let $\delta^{t}(f)$ be an interpolation function which interpolates between the scales $(j_{1},j_{2},\ldots,j_{d})$ of the averaging operators with scaling parameters $(t_{1},t_{2},\ldots,t_{d})$ with $0\leq t_{k}\leq 1$ for $k=1,2\ldots,m$:
(24) $\delta^{t}(f)\coloneq\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)$
such that lemma 4.3 holds
From the previous two lemmas and the fundamental theorem of calculus $A(f)$ has an identity which is a summation (with respect to the scaling parameters in all directions) over an integration of the power series comprised of the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d}$ with respect to the interpolation parameters in all directions. First, we define the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ which come from differentiating $A(\delta^{t}(f))$. The elements in $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ are defined such that each power of $\mathbf{v}$ corresponds to the number of independent partial derivatives of $\delta^{t}(f)$ taken:
(25) $\mathbf{v}^{1}(f)\coloneq\frac{\partial^{d}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{1}}(\delta^{t}(f))$
$\mathbf{v}^{2}(f)\coloneq\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\frac{\partial}{\partial t_{2}}(\delta^{t}(f))\frac{\partial}{\partial t_{1}}(\delta^{t}(f)))+$
(26) $\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{d-1}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{1}}(\delta^{t}(f))$
$\vdots$
$\mathbf{v}^{m}(f)\coloneq\frac{\partial^{m}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
$+\frac{\partial^{m-1}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{d-m+1}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-m}\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\ldots+$
(27) $\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{m}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{d-m-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-m-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
$\vdots$
$\mathbf{v}^{d-1}(f)\coloneq\frac{\partial^{2}}{\partial t_{d}\partial t_{d-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{d-1}\partial t_{d-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
(28) $+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))$
(29) $\mathbf{v}^{d}(f)\coloneq\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
Note that the first, $\mathbf{v}^{1}$, and last, $\mathbf{v}^{d}$, operators have one term while the operators are comprised of a series - each term in the series consisting of the number of partial derivatives corresponding to the power of the operator (i.e. the term $\mathbf{v}^{m}$ has $m$ partial derivatives for each term in its series). The identity for $A(f)$ can compactly be written as:
$A(f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}A^{\prime}(\delta^{t}(f))\mathbf{v}^{1}+A^{\prime\prime}(\delta^{t}(f))\mathbf{v}^{2}+\ldots+$
(30) $A^{m}(\delta^{t}(f))\mathbf{v}^{m}+\ldots+A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}+A^{d}(\delta^{t}(f))\mathbf{v}^{d}dt_{1}dt_{2}\ldots dt_{d}$
We can now define the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, by replacing the sequence $(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$ with $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ - each element in the latter sequence is comprised with a wavelet coefficient in the direction of the partial derivative for an element in the former sequence.
(31) $\mathbf{\tilde{v}}^{1}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
$\tilde{\mathbf{v}}^{2}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{2}}(f)Q^{j_{1}}(f)+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)+\ldots$
(32) $+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{1}}(f)$
$\vdots$
$\tilde{\mathbf{v}}^{m}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)$
$+Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{d-m+1}}(f)Q^{j_{d-m}}Q^{j_{m}}(f)\ldots Q^{j_{m-1}}(f)+\ldots+$
(33) $Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{d-m-1}}(f)Q^{j_{d-m-2}}(f)\ldots Q^{j_{1}}(f)$
$\vdots$
$\tilde{\mathbf{v}}^{d-1}(f)\coloneq Q^{j_{d}}Q^{j_{d-1}}(f)Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)+Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}(f)\ldots Q^{j_{1}}(f)$
(34) $+\ldots+Q^{j_{d}}(f)Q^{j_{d-1}}(f))\ldots Q^{j_{3}}(f)Q^{j_{2}}Q^{j_{1}}(f)$
(35) $\tilde{\mathbf{v}}^{d}(f)\coloneq Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)$
Again, note the first, $\tilde{\mathbf{v}}^{1}$, and last, $\tilde{\mathbf{v}}^{d}$, operators have one term while the operators, $(\mathbf{v}^{2},\ldots,\mathbf{v}^{d-1})$, are comprised of a series each with d1d-1 terms. In this case, each term in the series contains the number of wavelet coefficients corresponding to the power of the operator (i.e. the term $\tilde{\mathbf{v}}^{m}$ has $m$ wavelet coefficients for each term in its series). Now the following linear combination of $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ is independent of $t$ and is used to construct the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$:
$\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\coloneq\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}+A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2}+\ldots+A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m}$
$+\ldots+A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$
(36) $=\sum_{j_{d}=0_{d}}^{N_{d}}\sum_{j_{d-1}=0_{d-1}}^{N_{d-1}}\ldots\sum_{j_{1}=0_{1}}^{N_{1}}(\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta})$
Computing the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\coloneq A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, yields:
[…]
$(A^{\prime\prime}(\delta^{t}(f))\mathbf{v}^{2}-A^{\prime\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{2})+\ldots+(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})+\ldots+$
$(A^{d-1}(\delta^{t}(f))\mathbf{v}^{d-1}-A^{d-1}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d-1})+(A^{d}(\delta^{t}(f))\mathbf{v}^{d}-A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})dt_{1}dt_{2}\ldots dt_{d}$
(37) $=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}\int_{0}^{1}\int_{0}^{1}\ldots\int_{0}^{1}\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})dt_{1}dt_{2}\ldots dt_{d}$
We characterize the regularity of the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, by estimates on the terms in the integrand. Consider an arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the series, $\sum_{m=1}^{d}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand. Then
$\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
from Hölders inequality. Handling the estimates of $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ separately, we have:
$\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$
$=\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$
$\leq\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{1}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\int_{\alpha^{\mathbf{j}}_{\mathbf{k}}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}|A^{m}(y)|dy$
$\leq\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
(39) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
where $\alpha^{\mathbf{j}}_{\mathbf{k}}$ is the wavelet coefficient at the specified scale and location parameters, the second inequality comes from triangle inequality, the third inequality from the $L^{1}$ estimate on $A^{m}$- which is finite since $A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\in L^{\infty}$ by the maximum principle - and that wavelet coefficients of Hölder functions can be characterized by exponential decay in the scale parameters [4, 3]. For $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$,
$\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
$=(d-1)\lVert Q^{j_{1},j_{2},\ldots,j_{d}}\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
(40) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
where again, the last inequality comes from exponential decay in the wavelet scaling parameter and Hölder functions. Combining the estimates on $\lVert A^{m}(\delta^{t}(f))\rVert_{L^{1}([0,1]^{d})}$ and $\lVert\mathbf{v}^{m}\rVert_{L^{\infty}([0,1]^{d})}$ yields:
(41) $\lVert A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$
Now we bound the approximation terms, $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}$, for the mth term in the series under the integral for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ similarly. By Hölder’s inequality,
(42) $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}$
where
$\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}\leq\lVert A^{m}(Q^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$
(43) $\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
where the last inequality holds from the same argument in (39), and
$\lVert\tilde{\mathbf{v}}^{m}\rVert_{L^{\infty}([0,1]^{d})}=\lVert Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{m}}(f)Q^{j_{m-1}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}+$
$\lVert Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{d-m+1}}(f)Q^{j_{m-1}}(f)Q^{j_{m}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$+\ldots+\lVert Q^{j_{d}}(f)Q^{j_{d-1}}Q^{j_{d-2}}\ldots Q^{j_{d-m-1}}(f)Q^{j_{d-m-2}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{d}+j_{d-1}+\ldots+j_{m})\alpha}2^{-j_{m-1}\alpha}\ldots 2^{-j_{1}\alpha}$
$+2^{-(j_{d}+j_{d-1}+\ldots+j_{d-m+1})\alpha}2^{-j_{m-1}\alpha}\ldots 2^{-j_{1}\alpha}+\ldots+2^{-j_{d}\alpha}2^{-(j_{d-1}+\ldots+j_{d-m-1})\alpha}2^{-j_{d-m-2}\alpha}\ldots\ldots 2^{-j_{1}\alpha}$
(44) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
giving us,
(45) $\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
Combining the inequalities in (41) and (45) gives:
(46) $\left|\lVert(A^{m}(\delta^{t}(f))\mathbf{v}^{m}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
The argument above used to provide estimates for the arbitrary term, $(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$, in the integrand holds for the terms in the series, $\sum_{m=2}^{d-1}(A^{m}(\delta^{t}(f))\mathbf{v}^{m}-A^{m}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{m})$. The first and last terms are handled slightly different since $\mathbf{v}^{1}$ and $\mathbf{v}^{d}$ contain 1 term in their series instead of d1d-1 terms. Since $\mathbf{v}^{1}=\tilde{\mathbf{v}}^{1}$, the first term becomes:
(47) $(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}$
Then the following $L^{1}$ estimate is obtained:
$\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}}^{1}\rVert_{L^{1}([0,1]^{d})}\leq\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\rVert_{L^{1}([0,1]^{d})}\rVert\tilde{\mathbf{v}}^{1}\lVert_{L^{\infty}([0,1]^{d})}$
$=\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\lVert\delta^{t}(f))-P^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{1}([0,1]^{d})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{1}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\int_{\alpha^{\mathbf{j}}_{\mathbf{k}}(I^{j}_{k}\times I^{j}_{k}\times\ldots\times I^{j}_{k})}Q(y)dy\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
(48) $\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
by application of Hölder’s inequality, the property of $A\in C^{d}$, and the estimates on the wavelet coefficients of Hölder functions previously discussed. So by the previous result, the first term in the series under the integrand of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ has the estimate:
(49) $\left|\lVert(A^{\prime}(\delta^{t}(f))-A^{\prime}(P^{j_{1},j_{2},\ldots,j_{d}}(f)))\tilde{\mathbf{v}})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
Now we handle the final term in the series, $(A^{d}(\delta^{t}(f))\mathbf{v}^{d}-A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$, where $\tilde{\mathbf{v}}^{d}$ only has 1 term in its series. Since $\mathbf{v}\neq\tilde{\mathbf{v}}^{d}$, we handle $(A^{d}(\delta^{t}(f))\mathbf{v}^{d}$ and $A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$ separately as we did for the $m$th term; the obtained estimates are the same but the technology used is negligibly different. The $\lVert A^{d}(\delta^{t}(f))\rVert_{L^{1}([0,1])}$ estimate is the same as (40), but $\lVert\mathbf{v}^{d}\rVert_{L}^{\infty}([0,1])$ is slightly different:
$\lVert\mathbf{v}^{d}\rVert_{L^{\infty}([0,1]^{d})}\leq\lVert\sup_{t_{1},t_{2},\ldots,t_{d}}\mathbf{v}^{d}\rVert_{L^{\infty}([0,1]^{d})}$
$=\lVert Q^{j_{1},j_{2},\ldots,j_{d}}\rVert_{L^{\infty}([0,1]^{d})}$
$\leq C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\lVert Q^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f))\rVert_{L^{\infty}(I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}})}$
(50) $=C\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
Thus, we again have:
(51) $\lVert A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots,j_{d})2\alpha}$
as in (41). Similar reasoning applies to the $A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d}$ term. The $\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\rVert_{L^{1}([0,1]^{d})}$ estimate is the same as (43), but for $\lVert\tilde{\mathbf{v}}^{d}\rVert_{L^{\infty}([0,1]^{d})}$ we have:
$\lVert\tilde{\mathbf{v}}^{d}\rVert_{L^{\infty}([0,1]^{d})}=\lVert Q^{j_{d}}(f)Q^{j_{d-1}}(f)\ldots Q^{j_{1}}(f)\rVert_{L^{\infty}([0,1]^{d})}$
$\leq\sum_{k_{d}=1_{d}}^{2^{j_{d}}}Q^{j_{d}}_{k_{d}}(f)\sum_{k_{d-1}=1_{d-1}}^{2^{j_{d-1}}}Q^{j_{d-1}}_{k_{d-1}}(f)\ldots\sum_{k_{1}=1_{1}}^{2^{j_{1}}}Q^{j_{1}}_{k_{1}}(f)$
$\leq C\sum_{k_{d}=1_{d}}^{2^{j_{d}}}2^{-j_{d}\alpha}\sum_{k_{d-1}=1_{d-1}}^{2^{j_{d-1}}}2^{-j_{d-1}\alpha}\ldots\sum_{k_{1}=1_{1}}^{2^{j_{1}}}2^{-j_{1}\alpha}$
(52) $=C2^{-(j_{1}+j_{2}+\ldots+j_{d})\alpha}$
Combining (51) and (52), the last term of the series in the integrand of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ has the estimate:
(53) $\left|\lVert(A^{d}(\delta^{t}(f))\mathbf{v}^{d}\rVert_{L^{1}([0,1]^{d})}-\lVert A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{d})\rVert_{L^{1}([0,1]^{d})}\right|\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}$
So by (46), (49), and (53) we can express the residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, defined in (37), as a sum of piecewise constant functions:
$\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\sum_{j_{d}=0}^{N_{d}}\sum_{j_{d-1}=0}^{N_{d-1}}\ldots\sum_{j_{1}=0}^{N_{1}}(\sum_{k_{1}=1_{1}}^{2^{j_{1}}}\sum_{k_{2}=1_{2}}^{2^{j_{2}}}\ldots\sum_{k_{d}=1_{d}}^{2^{j_{d}}}\eta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
(54) $|\beta^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$
Now we characterize the regularity of $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ with the following lemma:
Lemma 4.5.
The residual, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, can be writen as:
$\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}(\sum_{k_{1},\ldots,k_{d}=1}^{2^{j_{1}},\ldots,2^{j_{d}}}\eta^{(j_{i})_{i=1}^{d}}_{(k_{i})_{i=1}^{d}}\chi_{I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}}(x_{1},x_{2},\ldots,x_{d})),$
(55) $|\eta^{(j_{i})_{i=1}^{d}}_{(k_{i})_{i=1}^{d}}|\leq C2^{-(j_{1}+j_{2}+\ldots j_{d})2\alpha}$
and is an element of $\Lambda_{2\alpha}([0,1]^{d})$
Proof.
To show $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\in\Lambda_{2\alpha}([0,1]^{d})$, we need to show
(56) $|\Delta_{(N_{i})_{i=1}^{d}}(A,f)(\mathbf{x})-\Delta_{(N_{i})_{i=1}^{d}}(A,f)(\mathbf{\tilde{x}})|\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
[…]
Similarly, let $\mathcal{L}$ be the set containing combinations of scaling and location parameters such that $\mathbf{x},\tilde{\mathbf{x}}$ are never contained in the same box.
[…]
Then summing over the scaling and location parameters in $\mathcal{U}$ for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$, one has:
$\sum_{(\mathbf{j},\mathbf{l})\in\mathcal{U}}\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{x})-\gamma^{\mathbf{{j}}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{\tilde{x}})=0$
(59) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
since $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ is piecewise constant on dyadic boxes (see 54). Now consider the case where $x,x^{\prime}$ are not contained in same dyadic box. Again, summing over the scaling and location parameters in $\mathcal{L}$ for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ one has:
$\sum_{(\mathbf{j},\mathbf{l})\in\mathcal{L}}\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{x})-\gamma^{\mathbf{j}}_{\mathbf{k}}\chi_{B^{\mathbf{j}}_{\mathbf{k}}}(\mathbf{\tilde{x}})\leq C\sum_{\mathbf{j}^{(i)}=\mathbf{j}^{(0)}}^{\mathbf{j}^{(N)}}2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(i)})2\alpha}$
$=\sum_{\mathbf{j}^{(i)}=\mathbf{j}^{(0)}}^{\mathbf{j}^{(N)}}2^{-(j^{(i)}_{1}+j^{(i)}_{2}+\ldots+j^{(i)}_{d})2\alpha}$
$\leq C2^{-(j^{(N)}_{1}+j^{(N)}_{2}+\ldots+j^{(N)}_{d})2\alpha}$
(60) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
where we order the scales in $\mathcal{L}$ such that $\mathbf{j}^{(0)}\leq\mathbf{j}^{(1)}\leq\ldots\leq\mathbf{j}^{(N)}$ indicates $2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(0)})}\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(1)})}\leq\ldots\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(N)})}$. The first inequality holds from the estimates on $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ obtained in (54), the second inequality from collapsing the geometric series, the third from the assumption on the scaling parameters contained in $\mathcal{L}$. Thus, combining the previous two estimates in (59) and (60) one has:
[…]
(61) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
Since the choice of $\mathbf{x},\mathbf{\tilde{x}}$ were arbitrarily chosen, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\in\Lambda_{2\alpha}([0,1]^{d})$
Thus, from lemma 4.5, we can compute the approximation $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ to $A(f)$ such that $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{d})$.
5. Paraproduct Caricatures
Included in this section are caricatures of d-tensor paraproducts stated in the main theorem.
Corollary 5.1.
Suppose $A\in C^{3}(\mathbb{R}),f\in\Lambda_{2\alpha}([0,1]^{3})$, then we obtain the following discrete 33-tensor paraproduct decomposition
$\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},j_{2},j_{3}=0}^{N_{1},N_{2},N_{3}}A^{\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{1}}(f)Q^{j_{2}}(f)Q^{j_{3}}(f)+$
$+A^{\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))[Q^{j_{2}}[Q^{j_{1}}(f)Q^{j_{3}}(f)]+Q^{j_{3}}[Q^{j_{2}}(f)Q^{j_{1}}(f)]]$
(62) $+A^{\prime\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{3}}Q^{j_{2}}Q^{j_{1}}(f)$
Proof.
The proof follows from the same techniques used to prove the general $d$-tensor case by setting $d=3$
Corollary 5.2.
The approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, is real analytic on the image of $f$ and has the following taylor expansion
(63) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$
where $A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ are the derivatives of A of order $\beta$ evaluated at the average of $f$ and $\tilde{\mathbf{v}}^{\beta}$ are wavelet operators acting in $\beta$ directions.
Proof.
As any real analytic function can be represented by its taylor series, it suffices to show AA is real analytic on $f$. Since $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ by Theorem 4.1 the power series representation $\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ is immediate, as $\beta$ represents the derivatives of $A$ and the powers of $\tilde{\mathbf{v}}$
6. Cone singularity in complex plane (d=3)
For the 3D example, we consider a time-varying singularity in the complex plane:
(64) $f(z)=\begin{cases}(1-|z|)^{\alpha}&\text{if }|z|<1\\ (1-\frac{1}{|z|})^{\alpha}&\text{if }\frac{1}{|z|}<1\end{cases}$
where $|z|=\sqrt{x^{2}+iy^{2}+t^{2}}$ is the usual modulus in the complex plane with a time-dependent parameter and $x,iy\in[-1,1],t\in[0,1]$. $N=256$ equispaced points are sampled in the x and y directions, 128 points are sampled in the time direction. Here, $A(f)$ is the same time-dependent graphic equalization function described in [2], where the amplitude of $f$ is differentially modulated at different frequency bands given by the scales, $j_{x}$ and $j_{iy}$, of the real and imaginary components:
(65) $A_{3}(f)=sin((\frac{j_{x}+j_{iy}}{2}<f(z),\psi^{j_{x},j_{iy}^{\prime}}(x,iy)>\psi^{j_{x},j_{iy}^{\prime}}(x,iy))$
Figure 1. Cone singularity in complex plane: (a - c) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=15 (d- f) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=70, (g-i) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=115.
The left column of Figure 1 shows the original singularity progressing towards the center, the middle column the nonlinear distortion, the right column the approximation. The rows represent discrete time points, $t=\{15,70,115\}$. Qualitatively, one can see the approximation recovers the ring singularity which is conspicuously distorted by $A(f)$.
We process all the figures in Figure 1 with the tensor haar basis. The red squares are the support of the tensor haar functions with the largest coefficients. This exemplifies the ability of the decomposition to recover the singularity in the presence of smooth non-linear distortions - the singularity is detected as expected for the original function, $f$, goes undetected for the non-linear distortion in the middle column, $A(f)$, and is recovered again by the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ in the right column.